In a geometric progression, the sum of the second and third term is 6 and the third and fourth term is -12, find the first, the common ratio and the sum of the first 10 terms of the progression


1
Expert's answer
2022-01-25T06:50:50-0500

The general term of a GP is:


"T_n = ar^{n-1}"

Therefore,


"T_2 = ar^{2-1} = ar\\
T_3 = ar^{3-1} = ar^2\\
T_4 = ar^{4-1} = ar^3"

Now,


"T_2+T_3 = 6 \implies ar+ar^2 =6 \qquad \qquad(i)\\
T_3+T_4 = -12 \implies ar^2+ar^3 =-12 \quad \quad(ii)"

From (ii):


"r(ar+ar^2) = -12 \qquad \qquad (iii)"

put (i) in (iii):


"6r =-12 \implies \textbf{r=-2}"

Put r=-2 in (i):


"a(-2) + a(-2)^2=6\\
-2a +4a=6\\
2a=6\\
\textbf{a=3}"

The sum of first n term of a GP is:


"S_{n}=a \cdot \frac{1-r^{n}}{1-r}\\
\text{ Given n=10, and putting other values to above formula: }\\

\begin{aligned}
S_{10} &=3 \cdot \frac{1-(-2)^{10}}{1-(-2)} \\
S_{10} &=3 \cdot(-341) \\
S_{10} &=-1023
\end{aligned}"